OpenStudy (anonymous):

17. Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (alekos):

have you attempted this?

OpenStudy (anonymous):

I am not good at algebra and yes i have

OpenStudy (alekos):

show me what you have done

OpenStudy (anonymous):

It is not that difficult. You are going to get everything on one side of the equation and then solve.

OpenStudy (alekos):

lets see how she goes

OpenStudy (anonymous):

8x-12=60 4x+24=60

OpenStudy (alekos):

where did the 60 come from?

OpenStudy (anonymous):

I would start by moving the x over to one side.

OpenStudy (anonymous):

If I understand the algebra I wouldnt be asking for help. I used a algebra calculator

OpenStudy (alekos):

oh i see...... you need to get the x's to the LHS and the constants to the RHS

OpenStudy (anonymous):

Algebra is a critical skill. We understand that you are asking for help, it is just that we don't want to give you the answer. It is important you learn the steps to get to the answer.

OpenStudy (anonymous):

I still dont know how to do that I dont want no one to give me that answer I just want examples how to solve them thats all. I aint asking for someone to give me that answers

OpenStudy (alekos):

in order to remove the 4x from the RHS what do you think we should do?

OpenStudy (anonymous):

Look I dont know

OpenStudy (anonymous):



OpenStudy (anonymous):

Sorry the drawing is not that great but it should get you started

OpenStudy (alekos):

i'll let jtryon assist you

OpenStudy (anonymous):

do i take away 8x from the 4x

OpenStudy (anonymous):

8x-4x=4x 12+24=36

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

You have 8x-12 = 4x+24. You are going to subtract 4x from both sides and then we have 4x-12 = 24.

OpenStudy (anonymous):

We want to get x by itself so we are going to add 12 to both sides and then solve for x.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

@MrsPurvis2010

OpenStudy (anonymous):

oh ok 8x-12+12=8x but how do i get 4x+24=12

OpenStudy (anonymous):

Ok let us start from the beginning. Your original equation is 8x - 12 = 4x + 24

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

When we subtract 4x from 4x, those two cancel out and 4x from 8x gives us 4x. So we have 4x - 12 = 24

OpenStudy (anonymous):

Do you follow so far?

OpenStudy (anonymous):

yea kinda

OpenStudy (anonymous):

do i - 8x-12-12 or do i add 8x-12+12

OpenStudy (anonymous):

You are not adding the whole 8x - 12... all that we are doing is subtracting 4x from the right side and 4x from the left side.

OpenStudy (anonymous):

4x-12 4x+12

OpenStudy (anonymous):

You just want to take it a step at a time.

OpenStudy (anonymous):

There you go. 4x - 12 = 24

OpenStudy (anonymous):

Correct that is what you have left

OpenStudy (anonymous):

yes please cause i am not really good at this

OpenStudy (anonymous):

now what do i do

OpenStudy (anonymous):

The next step is to get x on one side.

OpenStudy (anonymous):

Can you see how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

We have to isolate the variable so to do that, we want to add 12 to both sides.

OpenStudy (anonymous):

Your goal is to get x all by itself

OpenStudy (anonymous):

What happens when we add 12 to both sides?

OpenStudy (anonymous):

4x-12+12=4x 4x+12=12=4x+24

OpenStudy (anonymous):

i think you would divide 4x by 4

OpenStudy (anonymous):

@MrsPurvis2010

OpenStudy (anonymous):

Im still confused I might come back to it tomorrow

OpenStudy (anonymous):

That is the last step you do.

OpenStudy (anonymous):

When we add 12 to both sides, we end up with 4x = 36.

OpenStudy (butterflydreamer):